OpenStudy (h_hoff):

What conjecture can you make about the sum of the first 35 positive even numbers?

Parth (parthkohli):

Do you have to find the sum of the first 35 positive even numbers?

OpenStudy (h_hoff):

I have to find 2 numbers that multiply together to equal the sum of the first 35 positive numbers

Parth (parthkohli):

Ah, I see.

Parth (parthkohli):

Exercise: Find the sum of the first \(n\) positive even numbers. Hint: It is an arithmetic progression with first term \(2\), last term \(2n\) and number of terms \(n\).

OpenStudy (h_hoff):

Thank you

Parth (parthkohli):

Did you figure it out?

OpenStudy (anonymous):

\[2+4 + 6 +\cdots + 70=\\ 2(1+2+3+ +\cdots +35)=\\ \]

OpenStudy (anonymous):

Apply the formula \[1+2+3+\cdots+n=\frac{n(n+1)}2 \]

OpenStudy (anonymous):

\[1+2+3+\cdots +35=\frac {35\times 36}2\]

OpenStudy (anonymous):

\[2(1+2+3+\cdots+ 35)= 35\times 36 \]